# CREST: An Efficient Conjointly-trained Spike-driven Framework for Event-based Object Detection Exploiting Spatiotemporal Dynamics

### v1

@@ Acknowledgments
This work was supported by National Natural Science Foundation of China (No. 62074026) and NSAF under Grant U2030204. (Corresponding author: Jun Zhou.)
@@ B. Choice between ReLU and Leaky-ReLU
$U_{n}^{l}$ $=\sum_{m=1}^{N^{l-1}}\sum_{t=1}^{K^{l}}d^{l}(t)\beta^{l-1}\delta_{m}^{l-1}(t)w_{mn}^{l}$ (35)
$\delta_{n}^{l}(t)$ =$=\begin{cases}1,&u_{n}^{l}(t)\geq U_{th}^{l+1}(t)\\ -1,&u_{n}^{l}(t)\leq-U_{th}^{l+1}(t)\\ 0,&\text{otherwise }\end{cases}$ (36)
@@ C. Backpropagation Principles in DL-Net
$x^{l}_{n}=\sum_{m=1}^{N^{l-1}}x_{m}^{l-1}w_{mn}^{l}$ (39)
$x_{q}^{l}=\text{clip}(x^{l}_{n},0,X^{l})=\begin{cases}0,&x^{l}_{n}<0\\ x^{l}_{n},&0\leq x^{l}_{n}\leq X^{l}\\ X^{l},&\text{otherwise}\end{cases}$ (40)
$x_{p}^{l}=\text{round}(\frac{x_{q}^{l}}{X^{l}_{\text{min}}})\times X^{l}_{\text{min}}$ (41)
@@
Then, from Eq. 9,10,11 , the gradient of $w_{mn}^{L}$ can be obtained as:
$\frac{\partial\mathcal{L}}{\partial w_{mn}^{L}}=\frac{\partial\mathcal{L}}{\partial x_{n}^{L}}\frac{\partial x_{n}^{L}}{\partial w_{mn}^{L}}=\epsilon_{n}^{L}\frac{\partial x_{p}^{L-1}w_{mn}^{L}}{\partial w_{mn}^{L}}=\epsilon_{n}^{L}x_{p}^{L-1}$ (44)
@@
$\quad=\sum_{n=1}^{N^{L}}w_{mn}^{L}\frac{\partial x_{q}^{L-1}}{\partial x_{m}^{L-1}}\epsilon_{n}^{L}$ (45)
@@ D. Datasets details
NCAR dataset.
The NCAR dataset (Sironi et al. 2018) are choosed for event-based object recognition task, which comprising 12,336 car samples and 11,693 background samples. Each sample has a duration of 100 ms and exhibits varying spatial dimensions.

### v3

@@ Acknowledgments
This work was supported by NSAF under Grant U2030204. (Corresponding author: Jun Zhou.)
@@ B. Choice between ReLU and Leaky-ReLU
$U_{n}^{l}$ $=\sum_{m=1}^{N^{l-1}}\sum_{t=1}^{K^{l-1}}d^{l-1}(t)\beta^{l-1}\delta_{m}^{l-1}(t)w_{mn}^{l}$ (35)
$\delta_{n}^{l}(t)$ =$=\begin{cases}1,&u_{n}^{l}(t)\geq U_{th}^{l}(t)\\ -1,&u_{n}^{l}(t)\leq-U_{th}^{l}(t)\\ 0,&\text{otherwise }\end{cases}$ (36)
@@ C. Backpropagation Principles in DL-Net
$x_{q}^{l-1}=\text{clip}(x^{l-1}_{m})=\begin{cases}0,&x^{l-1}_{m}<0\\ x^{l-1}_{m},&0\leq x^{l-1}_{m}\leq X^{l-1}\\ X^{l-1},&\text{otherwise}\end{cases}$ (39)
$x_{p}^{l-1}=\text{round}(\frac{x_{q}^{l-1}}{X^{l-1}_{\text{min}}})\times X^{l-1}_{\text{min}}$ (40)
$x^{l}_{n}=\sum_{p=1}^{N^{l-1}}x_{p}^{l-1}w_{pn}^{l}$ (41)
@@
Then, from Eq. 39,40,41, the gradient of $w_{pn}^{L}$ can be obtained as:
$\frac{\partial\mathcal{L}}{\partial w_{pn}^{L}}=\frac{\partial\mathcal{L}}{\partial x_{n}^{L}}\frac{\partial x_{n}^{L}}{\partial w_{pn}^{L}}=\epsilon_{n}^{L}\frac{\partial x_{p}^{L-1}w_{pn}^{L}}{\partial w_{pn}^{L}}=\epsilon_{n}^{L}x_{p}^{L-1}$ (44)
@@
$\quad=\sum_{n=1}^{N^{L}}w_{pn}^{L}\frac{\partial x_{q}^{L-1}}{\partial x_{m}^{L-1}}\epsilon_{n}^{L}$ (45)
@@ D. Datasets details
NCARs dataset.
The NCARs dataset (Sironi et al. 2018) are choosed for event-based object recognition task, which comprising 12,336 car samples and 11,693 background samples. Each sample has a duration of 100 ms and exhibits varying spatial dimensions.